# From Local to Global Order: A Theory of Neural Synaptic Balance

### v2

@@ 12. Conclusion
The theory of neural synaptic balance explains some basic findings regarding $L_{2}$ balance in feedforward networks of ReLU neurons and extends them in several directions. The first direction is the extension to BiLU and other activation functions (BiPU). The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied in a stochastic or deterministic manner, global order always emerges through the convergence of the balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. Scaling and balancing operations are local and thus may have applications in physical, non-digitally simulated, neural networks where the emergence of global order from local operations may lead to better operating characteristics and lower energy consumption.
@@
Thus, in short, we obtain a unique closed-form expression for each $M_{i}$. From there, we infer the unique and final state of the weights, where $A^{*}_{i}=M_{i}A_{i}=\Lambda_{i}A_{l}/\Lambda_{l-1}$. Note that each $M_{i}$ depends on all the other $M_{j}$’s, again showcasing how the local balancing algorithm leads to a unique global solution. ∎

### v3

@@ 13. Conclusion
The theory of synaptic neural balance explains some basic findings regarding $L_{2}$ balance in feedforward networks of ReLU neurons and extends them in several directions. The first direction is the extension to BiLU and other activation functions (BiPU). The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied in a stochastic or deterministic manner, global order always emerges through the convergence of the balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. Balancing can be applied, fully or partially, before, during, or at after training and can help improve the convergence or final performance. Scaling and balancing operations are local and thus may have applications in physical, non-digitally simulated, neural networks where the emergence of global order from local operations may lead to better operating characteristics and lower energy consumption.
Contributions: The theory was developed by PB. The simulations to corroborate the theory were carried out by A. R. The simulations to improve training were carried out by I.D. and A. A. The manuscript was written by P.B with input from all the other authors. All authors reviewed the final version.
@@
Thus, in short, we obtain a unique closed-form expression for each $M_{i}$. From there, we infer the unique and final state of the weights, where $A^{*}_{i}=M_{i}A_{i}$. Note that each $M_{i}$ depends on all the other $M_{j}$’s, again showcasing how the local balancing algorithm leads to a unique global solution. ∎